OpenStudy (anonymous):

Can somebody tell me how to find the anti-derivative of cos(x) * x and cos(x) * sin(x)

jimthompson5910 (jim_thompson5910):

to integrate cos(x) * x, I'd use the Integration by Parts rule http://tutorial.math.lamar.edu/Classes/CalcII/IntegrationByParts.aspx

jimthompson5910 (jim_thompson5910):

to integrate cos(x) * sin(x), use the identity sin(2x) = 2*sin(x)*cos(x)

jimthompson5910 (jim_thompson5910):

then use u-substitution

OpenStudy (kainui):

The first one requires a clever use of the product rule. The second one requires a clever use of the chain rule. product rule (f*g)'=f'g+fg' chain rule (f(g))'=f'(g)*g'